OpenStudy (anonymous):

sin(3pheta/2)(cospheta/2) express each product as a sum containing only sines or cosines. step by step

OpenStudy (anonymous):

Its not coming out right. i tried.

OpenStudy (dumbcow):

\[\sin(3x/2) \cos(x/2) = \frac{\sin (3x/2 + x/2) + \sin(3x/2 - x/2)}{2}\]

OpenStudy (anonymous):

okay, thank you.